# SOFIA/HAWC+ Far-Infrared Polarimetric Large Area CMZ Exploration Survey. V. The Magnetic Field Strength and Morphology in the Sagittarius C Complex

### v1

@@ 3.2.1 Angular Dispersion ($\sigma_{\phi}$)
Past decades of research have shown that the GC magnetic field exhibits different behaviours viewed from different observational scales: a pervasive and ordered magnetic field in large-scale (Novak et al., 2003; Planck Collaboration et al., 2015; Mangilli et al., 2019; Guan et al., 2021), a turbulent magnetic field driven by local environments in molecular clouds in small-scale (Chuss et al., 2003; Butterfield et al., 2024a, b; Lu et al., 2024; Paré et al., 2025), and a vertical inter-cloud field signified by the NTFs (Serabyn & Morris, 1994). However, in a dynamically complex region like the CMZ, the relative contribution of magnetic field systems on different scales is still under investigation. To this end, we follow the methods developed by Hildebrand et al. (2009) and Houde et al. (2009), which have been implemented in numerous SOFIA/HAWC+ studies (Chuss et al., 2019; Guerra et al., 2021; Butterfield et al., 2024a, b). We direct interested readers to Houde et al. (2009, 2016) for a detailed account of this method, and note that the analysis in this subsection was implemented using the polBpy library (Guerra et al., 2024).
@@
where $\bar{P}({\bf r})$ is the beam-integrated polarized flux at coordinate ${\bf r}$. Instead of using the default array-by-array approach taken by the polBpy package, we use a Fast Fourier Transform to compute $\langle\bar{P}^{2}(l)\rangle$. Intuitively, Equation 6 is a statistical measure of the average polarized flux within a distance $l$ in the sky plane, thereby giving the approximate spatial extent of a cloud’s projection in the sky-plane. Assuming the cloud is isotropic, we posit that the distribution of polarized flux along the cloud’s depth is similar. We follow Houde et al. (2009) and use the half-width at half-maximum (HWHM) value as the effective cloud depth, as it serves as a good approximation of the spatial extent of a cloud that contains the bulk of the polarized flux. However, because the Sgr C complex consists of several regions identifiable by their spatial, radial velocity separation, or by the overall trends in their magnetic field orientations, the dispersion calculation only remains meaningful if applied to individual regions with their own identifiable magnetic field system. The regions that we have thereby identified, along with their spatial extents, are shown in Figure 4.
@@
Having estimated the effective cloud depth, we perform a Markov Chain Monte Carlo (MCMC) fit of Equation 4 for variables $\delta$, $a_{2}$, and most importantly, $\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$ for all pixels in the field of view. The polBpy package uses the emcee package (Foreman-Mackey et al., 2013) to generate MCMC posterior distributions of parameters $\delta$ and $a_{2}$, but do so for $\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$ through a similar quantity $\Delta^{\prime}(\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle)^{-1}$. All parameters are constrained to be positive with no upper bounds, though we do note that $\delta$ cannot be greater than the diameter of the dispersion function’s circular kernel. The fit was performed on the structure function with $l\leq 1.25$′ to ensure only local dispersion is sampled. Furthermore, to improve the efficiency of the MCMC fits, we calculate and fit the dispersion functions using all available polarization measurements within one region shown in Figure 4. The best-fit results are then used as the initial condition of the pixel-by-pixel MCMC fit to accelerate convergence. The final results were obtained with 300 iterations for each pixel, at which point we manually inspected the MCMC fit at numerous locations to ensure convergence. We found that the fit results do not evolve significantly after 150 iterations. The resulting images of the variables are shown in Figure 5.
@@
Globally, the $a_{2}$ and $\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$ values are positively correlated, and they are negatively correlated with $\delta$. Given the high column density in many clouds in Sgr C, an increase in turbulence (which entails a decrease in turbulent correlation length) likely decreases the polarization fraction integrated along a line of sight. Therefore, the large-scale contribution ${\bf B}_{0}$ naturally becomes the dominant term relative to the turbulent contribution ${\bf B}_{t}$ in highly turbulent regions.
@@ 3.2.2 Mass Density ($\rho_{m}$)
We infer the mass density by deriving a column density image using the Hi-GAL survey data (Molinari et al., 2010, 2016). Our procedure largely follows the spectral energy distribution (SED) fitting process described in section 3.1.3 of Chuss et al. (2019) and section 2.2 of Paré et al. (2025). Our SED fitting includes the Hi-GAL survey images from 160 µm, 250 µm, and 350 µm. We did not include the 70 µm image due to its uncertain extinction correction (Compiègne et al., 2010; Battersby et al., 2011, 2024). The 500 µm was also excluded due to its low resolution. The images are convolved with a Gaussian kernel to match the 24″ resolution of the SPIRE 350 µm image. Finally, we perform a pixel-by-pixel SED fitting to the single-temperature modified blackbody modelled by Vaillancourt (2002)
@@
where $\kappa_{\nu_{0}}$ is the reference dust opacity, $\mu$ is the mean molecular weight per hydrogen atom, and $m_{\rm H}$ is the atomic mass of hydrogen. We adopt ${\nu_{0}}=1000~{}{\rm GHz}$, $\kappa_{\nu_{0}}(1000~{}{\rm GHz})=0.1~{}\mathrm{cm^{2}\,g^{-1}}$ and $\mu=2.8$. The derived column density map is shown in Figure 6.
Figure 6: The column density map derived from fitting Equation 7 to the Hi-GAL 160, 250, and 350 µm images. The blue circle and rectangle indicate the location of the Sgr C H II region and the Sgr C NTF.
As expected, the cold molecular clouds, such as FIR-4, the $-65$ km s-1 cloud, and the $-90$ km s-1 cloud, which dominates the far-IR emission, are the most prominent in this map. We compare our column density result with that of Battersby et al. (2024), which uses a slightly different method and also excludes the 70 µm image, as well as Paré et al. (2025), which used the same method with this work but included the 70 µm Hi-GAL image. The column density values are similar across these works upon comparison.
@@ 3.2.3 Velocity Dispersion ($\sigma_{v}$)
Due to the turbulent nature of Sgr C, we adopt a constant velocity dispersion of $12$ km s-1 by calculating the second spectral moment map of the HCN spectral cube provided by Jones et al. (2012) using the Common Astronomy Software Applications (CASA; CASA Team et al. 2022) package and obtain the dispersion by taking its square root.
@@ 3.2.4 Magnetic Field Strength Results
The sky-plane component of the magnetic field strength calculated from the above-described inputs using Equation 5 is displayed in the lower right panel of Figure 5. The field strength exhibits a large range from $\sim 30~{}\mu{\rm G}$ to $\sim 400~{}{\rm\mu G}$. This range agrees with many past attempts to infer the CMZ magnetic field strength (Crocker et al., 2010; Chuss et al., 2003; Kruijssen et al., 2014; Sormani et al., 2020; Tress et al., 2024). We find that the morphology of the angular or velocity dispersion image has a limited effect on the inferred $|{\bf B}_{\rm POS}|$, but the local turbulence of a region is anti-correlated with $|{\bf B}_{\rm POS}|$ (shown in the bottom panels of Figure 5). Therefore, the dominant factor affecting $|{\bf B}_{\rm POS}|$ is the local shear and compression. On the scale of individual molecular clouds, this conclusion suggests that the most turbulent clouds (FIR-4, the $-90$ km s-1 cloud, and the $-65$ km s-1 cloud) exhibit the lowest $|{\bf B}_{\rm POS}|\sim 100~{}\mu{\rm G}$.
@@
We note a few caveats in our DCF analysis. One should be cautious that the FIREPLACE survey is conducted in a single wavelength. The 214 µm filter tends to sample clouds with colder temperatures (colder compared to the warmer Sgr C H II region; Carey et al. 2009, Hankins et al. 2020), which in some regions may only be one of multiple clouds along the line of sight. The most obvious example in Sgr C is a cloud centered at $l=359.48$, $b=-0.17$ with a broad range of velocities (from -92.6 km s-1 to -34.3 km s-1) detected in the HCN line analysis presented in Section 4.2 (see Figure 8). The 850 µm measurements from Lu et al. (2024) nicely trace the morphology of this cloud but disagree with the 214 µm magnetic field orientations in this region, which may indicate that the 850 µm measurements are sampling a different cloud along that line of sight, which may naturally possess a different magnetic field orientation. As such, inferring $|{\bf B}_{\rm POS}|$ using only one wavelength introduces some uncertainty, since different clouds, or even the same cloud with different temperature domains, along a given line of sight can have independent field directions that then lead to a reduction of the polarization fraction. Another possible source of error comes from the column density estimation, as we assume a single-temperature blackbody, having multiple clouds (possibly with different temperatures) along the same line of sight could cause an error in our column density estimation, which also affects the $|{\bf B}_{\rm POS}|$ calculation. In regions dominated by higher temperatures, the DCF analysis cannot be reliably performed since there is an absence of valid 214 µm magnetic field detections.
@@ 4.1 Assembly of Observations
Figure 7: Selected multi-wavelength images of the Sgr C complex. The region has been surveyed in $1-4~{}{\rm keV}$ X-rays by Chandra (Wang, 2021), 8 µm by Spitzer/IRAC (Stolovy et al., 2006), 24 µm by Spitzer/MIPSGAL (Carey et al., 2009), 70 µm by Herschel/PACS (Molinari et al., 2010), 214 µm by SOFIA/HAWC+ (Paré et al., 2024), 450 µm and 850 µm by JCMT/SCUBA (Pierce-Price et al., 2000), 2 mm by IRAM/GISMO (Arendt et al., 2019), and at 20 cm by MeerKAT (Heywood et al., 2022). For visual comparison between the multi-wavelength images and the magnetic field detections, we add the 214 µm magnetic field pseudovectors from Paré et al. (2024) in Panel #5 as yellow segments. The corresponding beam sizes are displayed as white circles in the lower right corner of each image. All images are displayed with a logarithmic stretch.
@@
The Galactic plane appears bright in the IR regime out to 70 µm and in radio, with the peak emission at these wavelengths arising near the Sgr C H II region. In mid-IR and radio images (Panels #2–4, and #7 in Figure 7), the Sgr C H II region exhibits a similar elliptical morphology elongated along the Galactic plane. We observe an absorption “tail” originating from inside the Sgr C H II region pointing towards the southeast, where the EGO is located (G359.44-0.10; Kendrew et al. 2013). This low-intensity “valley” is observed at 8, 24, and 70 µm and is a result of foreground absorption by the head-tail cloud, which turns to emission in far-IR wavelengths beyond 70 µm. The mid-IR intensity images tend to become more diffuse at longer wavelengths, which is partially due to the trend that lower-temperature dust is generally the most extended, and partially due to decreased resolution at longer wavelengths.
@@
In the far-IR regime (214–850 µm, Panels #5-7 in Figure 7), the images are dominated by thermal emission from cold dust residing in molecular clouds. The Galactic plane and the H II region, by contrast, appear dark in the far-IR regime. The “head” of the head-tail structure is larger in size at far-IR wavelengths than in the mid-IR. Though this decrease in size could be a result of cold dust being more diffuse, it can again be a result of lower spatial resolution in far-IR wavelengths. The “tail” also appears bright at these wavelengths, reflecting a relatively high column density of cold dust. Indeed, we have an inferred column density of $\sim 10^{23}~{}\mathrm{cm^{-2}}$ in this region, shown in Figure 6.
A noteworthy structure is the vertical “stripe” extending northwards from $l=359.52$, $b=-0.05$ that appears in the 450 and 850 µm SCUBA images (Panels #6–7 in Figure 7), though it is most noticeable at 450 µm and is much more diffuse at 850 µm. With the caveat that this could be an artifact due to filtering, we note that the northern end of this “stripe” coincides with the mid-IR ridge that connects the extended thermal source AFGL5376 and the Galactic plane (Uchida et al., 1990, 1994). The 214 µm magnetic field orientations do not bear any obvious relationship to the morphology of this stripe. We go no further than noting the above correspondence, as a discussion of AFGL5376 is beyond the scope of this paper.

### v2

@@ 3.2.1 Angular Dispersion ($\sigma_{\phi}$)
Past decades of research have shown that the GC magnetic field exhibits different behaviours viewed from different observational scales: a pervasive and ordered magnetic field in large-scale (Novak et al., 2003; Planck Collaboration et al., 2015; Mangilli et al., 2019; Guan et al., 2021), a turbulent magnetic field driven by local environments in molecular clouds in small-scale (Chuss et al., 2003; Butterfield et al., 2024a, b; Lu et al., 2024; Paré et al., 2025), and a vertical inter-cloud field signified by the NTFs (Serabyn & Morris, 1994). However, in a dynamically complex region like the CMZ, the relative contribution of magnetic field systems on different scales is still under investigation. To this end, we follow the methods developed by Hildebrand et al. (2009) and Houde et al. (2009), which have been implemented in numerous SOFIA/HAWC+ studies (Chuss et al., 2019; Guerra et al., 2021; Butterfield et al., 2024a, b). We direct interested readers to Houde et al. (2009, 2016) for a detailed account of this method. The analysis in this subsection was implemented using the polBpy software (Guerra et al., 2024).
@@
where $\bar{P}({\bf r})$ is the beam-integrated polarized flux at coordinate ${\bf r}$. Instead of using the default array-by-array approach taken by the polBpy package, we use a Fast Fourier Transform to compute $\langle\bar{P}^{2}(l)\rangle$. Intuitively, Equation 6 is a statistical measure of the average polarized flux within a distance $l$ in the sky plane, thereby giving the approximate spatial extent of the part of a cloud having the bulk of a coherent magnetic field orientation in the sky-plane. Assuming the cloud is isotropic, we posit that the distribution of polarized flux along the cloud’s depth is similar to its sky-plane projection. We thereby follow Houde et al. (2009) and use the half-width at half-maximum (HWHM) value as the effective cloud depth. However, because the Sgr C complex consists of several regions identifiable by their spatial and radial velocity separation, or by the overall trends in their magnetic field orientations, the dispersion calculation only remains meaningful if applied to individual regions with their own identifiable magnetic field system. Therefore, we visually group the Sgr C magnetic fields based on their large-scale trends and only conduct DCF analyses using the magnetic field pseudovectors within a given region. The regions that we have thereby identified, along with their spatial extents, are shown in Figure 4.
@@
Having estimated the effective cloud depth, we perform a Markov Chain Monte Carlo (MCMC) fit of Equation 4 for variables $\delta$, $a_{2}$, and most importantly, $\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$ for all pixels in the field of view. The polBpy package uses the emcee package (Foreman-Mackey et al., 2013) to generate MCMC posterior distributions of parameters $\delta$ and $a_{2}$, but do so for $\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$ through a similar quantity $\Delta^{\prime}(\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle)^{-1}$. All parameters are constrained to be positive with no upper bounds, though we do note that $\delta$ cannot be greater than the diameter of the dispersion function’s circular kernel. The fit was performed on the structure function with $l\leq 1.25$′ to ensure that only local dispersion is sampled. Here, we note that the behaviour of the kernel on the boundaries between regions can lead to improperly defined dispersion functions and therefore less-then-optimal fitted parameters. However, the MCMC fitting routine solves for the quantity $\Delta^{\prime}(\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle)^{-1}$ for each pixel, which allows for using a pixel-dependent value of effective depth (see Guerra et al., 2021). Furthermore, to improve the efficiency of the MCMC fits, we calculate and fit the dispersion functions using all available polarization measurements within one region shown in Figure 4. The best-fit results are then used as the initial condition of the pixel-by-pixel MCMC fit to accelerate convergence. The final results were obtained with 300 iterations for each pixel, at which point we manually inspected the MCMC fit at numerous locations to ensure convergence. We found that the fit results do not evolve significantly after 150 iterations. The resulting images of the variables are shown in Figure 5.
@@
Globally, the $a_{2}$ and $\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$ values are positively correlated, and they are negatively correlated with $\delta$. Given the high column density in many clouds in Sgr C, an increase in turbulence (which entails a decrease in turbulent correlation length) likely decreases the polarization fraction integrated along a line of sight. Therefore, the large-scale contribution ${\bf B}_{0}$ naturally becomes the dominant term relative to the turbulent contribution ${\bf B}_{t}$ in highly turbulent regions. We also note that, given the issue of moving kernel near regional boundaries, the $\langle{\bf B}_{t}^{2}\rangle/\langle{\bf B}_{0}^{2}\rangle$ results on the boundaries should be taken with caution. The observed discontinuities in Figure 5 are more likely due to unresolved $\delta$ parameter in some regions, which is a consequence of limited angular resolution.
@@ 3.2.2 Mass Density ($\rho_{m}$)
We infer the mass density by deriving a column density image using the Hi-GAL survey data (Molinari et al., 2010, 2016). Our procedure largely follows the spectral energy distribution (SED) fitting process described in section 3.1.3 of Chuss et al. (2019) and section 2.2 of Paré et al. (2025). Our SED fitting includes the Hi-GAL survey images from 160, 250, and 350 µm. We did not include the 70 µm image due to its uncertain extinction correction (Compiègne et al., 2010; Battersby et al., 2011, 2024). The 500 µm image was also excluded due to its low resolution. The images are convolved with a Gaussian kernel to match the 24″ resolution of the SPIRE 350 µm image. Following the convolution, we perform a pixel-by-pixel SED fitting to the single-temperature modified blackbody modelled by Vaillancourt (2002)
@@
where $\kappa_{\nu_{0}}$ is the reference dust opacity, $\mu$ is the mean molecular weight per hydrogen atom, and $m_{\rm H}$ is the atomic mass of hydrogen. We adopt ${\nu_{0}}=1000~{}{\rm GHz}$, $\kappa_{\nu_{0}}(1000~{}{\rm GHz})=0.1~{}\mathrm{cm^{2}\,g^{-1}}$ and $\mu=2.8$ as the mean molecular weight of hydrogen. The derived column density map is shown in Figure 6.
Figure 6: The column density map derived from fitting Equation 7 to the Hi-GAL 160, 250, and 350 µm images, displayed in a logarithmic scale. The blue circle and rectangle indicate the locations of the Sgr C H II region and the Sgr C NTF, respectively.
As expected, the cold molecular clouds, such as FIR-4, the $-65$ km s-1 cloud, and the $-90$ km s-1 cloud, which dominate the far-IR emission, are the most prominent in this map. We compare our column density result with that of Battersby et al. (2024), which uses a slightly different method and also excludes the 70 µm image, as well as Paré et al. (2025), which used the same method as this work but includes the 70 µm Hi-GAL image. The column density values are similar across these works upon comparison. The mass density is then calculated in each region as
$\rho_{m}=\mu m_{H}\frac{N(H_{2})}{\Delta^{\prime}},$ (10)
following Guerra et al. (2024).
@@ 3.2.3 Velocity Dispersion ($\sigma_{v}$)
Due to the highly turbulent nature of the clouds surrounding Sgr C, the velocity dispersion is quite large, as in most of the CMZ. We measure the velocity dispersion by taking the square root of the second spectral moment map of the HCN spectral cube provided by Jones et al. (2012) using the Common Astronomy Software Applications (CASA; CASA Team et al. 2022) package. Over the entire region, the velocity dispersion varies by only about 20% around a mean value of $10$ km s-1, and the mean value within each of our defined regions varies by even less. Meanwhile, there are many instrumental artifacts contained in the velocity dispersion map calculated from Mopra spectral cube, which cannot be easily mitigated. Given the inherent uncertainty in our sky-plane magnetic field strength determinations owing to the fact that we only have access to the intensity-weighted average of the velocity dispersion along the line of sight, and similarly for the magnetic field orientations, we have chosen to adopt a constant velocity dispersion of $10$ km s-1. The resulting field strength estimates vary linearly with the adopted velocity dispersion (c.f., Equation 2).
@@ 3.2.4 Magnetic Field Strength Results
The sky-plane component of the magnetic field strength calculated from the above-described inputs using Equation 4 is displayed in the lower right panel of Figure 5. The field strength exhibits a large range from $\sim 30~{}\mu{\rm G}$ to $\sim 300~{}{\rm\mu G}$. This range agrees with many past attempts to infer the CMZ magnetic field strength (Crocker et al., 2010; Chuss et al., 2003; Kruijssen et al., 2014; Sormani et al., 2020; Tress et al., 2024). We find that the morphology of the angular or velocity dispersion image has a limited effect on the inferred $|{\bf B}_{\rm POS}|$, but the local turbulence of a region is anti-correlated with $|{\bf B}_{\rm POS}|$ (shown in the bottom panels of Figure 5). Therefore, the dominant factor affecting $|{\bf B}_{\rm POS}|$ is the local shear and compression. On the scale of individual molecular clouds, this conclusion suggests that the most turbulent clouds (FIR-4, the $-90$ km s-1 cloud, and the $-65$ km s-1 cloud) exhibit the lowest $|{\bf B}_{\rm POS}|\sim 100~{}\mu{\rm G}$.
@@
We note a few caveats in our DCF analysis. One should be cautious that the FIREPLACE survey is conducted in a single wavelength. The 214 µm filter tends to sample clouds with colder temperatures (colder compared to the warmer Sgr C H II region; Carey et al. 2009, Hankins et al. 2020), which in some regions may only be one of multiple clouds along the line of sight. The most obvious example in Sgr C is a cloud centred at $l=359.48$, $b=-0.17$ with a broad range of velocities (from -92.6 km s-1 to -34.3 km s-1) detected in the HCN line analysis presented in Section 4.2 (see Figure 8). The 850 µm measurements from Lu et al. (2024) nicely trace the morphology of this cloud but disagree with the 214 µm magnetic field orientations in this region, which may indicate that the 850 µm measurements are sampling a different cloud along that line of sight, which may naturally possess a different magnetic field orientation. As such, inferring $|{\bf B}_{\rm POS}|$ using only one wavelength introduces some uncertainty, since different clouds, or even the same cloud with different temperature domains, along a given line of sight can have independent field directions that then lead to a reduction of the polarization fraction. Another possible source of error comes from the column density estimation, as we assume a single-temperature blackbody, having multiple clouds (possibly with different temperatures) along the same line of sight could cause an error in our column density estimation, which also affects the $|{\bf B}_{\rm POS}|$ calculation. In regions dominated by higher temperatures, the DCF analysis cannot be reliably performed since there is an absence of valid 214 µm magnetic field detections.
@@ 4.1 Assembly of Observations
Figure 7: Selected multi-wavelength images of the Sgr C complex. The region has been surveyed in $1-4~{}{\rm keV}$ X-rays by Chandra (Wang, 2021), 8 µm by Spitzer/IRAC (Stolovy et al., 2006), 24 µm by Spitzer/MIPSGAL (Carey et al., 2009), 70 µm by Herschel/PACS (Molinari et al., 2010), 214 µm by SOFIA/HAWC+ (Paré et al., 2024), 450 µm and 850 µm by JCMT/SCUBA (Pierce-Price et al., 2000), 2 mm by IRAM/GISMO (Arendt et al., 2019), and at 20 cm by MeerKAT (Heywood et al., 2022). For visual comparison between the multi-wavelength images and the magnetic field detections, we add the 214 µm magnetic field pseudovectors from Paré et al. (2024) in Panel #5 as yellow segments. The corresponding beam sizes are displayed as white circles in the lower right corner of each image. The Green star in Panel #2 labels the G359.44-0.10 EGO, and the yellow arrows indicate in Panels #6 and #7 indicates the location of a vertical structure in these maps. All images are displayed with a logarithmic stretch.
@@
The Galactic plane appears bright in the IR regime out to 70 µm and in radio, with the peak emission at these wavelengths arising near the Sgr C H II region. In mid-IR and radio images (Panels #2–4, and #7 in Figure 7), the Sgr C H II region exhibits a similar elliptical morphology elongated along the Galactic plane. We observe an absorption “tail” originating from inside the Sgr C H II region extending towards the southeast, where the EGO is located (G359.44-0.10; Kendrew et al. 2013; Lu et al. 2019, 2021; Crowe et al. 2024; labelled in Panel #2 of Figure 7). This low-intensity “valley” is manifested at 8, 24, and 70 µm and is a result of foreground absorption by the head-tail cloud, which turns to emission at far-IR wavelengths beyond 70 µm. The mid-IR intensity images tend to become more diffuse at longer wavelengths, which is partially due to the trend that lower-temperature dust is generally the most extended, and partially due to decreased resolution at longer wavelengths.
@@
In the far-IR regime (214–850 µm, Panels #5–7 in Figure 7), the images are dominated by thermal emission from cold dust residing in molecular clouds. The Galactic plane and the H II region, by contrast, appear dark in the far-IR regime. The “head” of the head-tail structure is larger in size at far-IR wavelengths than in the mid-IR. Though this increase in size could be a result of cold dust being more diffuse, it can again be a result of lower spatial resolution at far-IR wavelengths. The “tail” also appears bright at these wavelengths, reflecting a relatively high column density of cold dust. Indeed, we have an inferred column density of $\sim 10^{23}~{}\mathrm{cm^{-2}}$ in this region, shown in Figure 6.
A noteworthy structure is the vertical “stripe” extending northwards from $l=359.52$, $b=-0.05$ that appears in the 450 and 850 µm SCUBA images ( labelled by orange arrows in Panels #6–7 in Figure 7), though it is most noticeable at 450 µm and is much more diffuse at 850 µm. With the caveat that this could be an artifact due to filtering, we note that the northern end of this “stripe” coincides with the mid-IR ridge that connects the extended thermal source AFGL5376 and the Galactic plane (Uchida et al., 1990, 1994). The 214 µm magnetic field orientations do not bear any obvious relationship to the morphology of this stripe. We go no further than noting the above correspondence, as a discussion of AFGL5376 is beyond the scope of this paper.